OpenStudy (anonymous):

What is the 7th term of the geometric sequence where a1 = -2 and a5 = -512?

OpenStudy (lgbasallote):

find the common ratio using the formula \[\Large A_n = A_1 r^{n-1}\] let \(\large A_n = A_5\) \[\Large A_5 = A_1r^{5-1}\] i've done enought..you do the rest :D

OpenStudy (anonymous):

ok thanks

OpenStudy (lgbasallote):

you know what to do after you get the r right?

OpenStudy (anonymous):

ill try

OpenStudy (lgbasallote):

okay :0

OpenStudy (anonymous):

:)

OpenStudy (lgbasallote):

yeah that was supposed to be a :) too lol

OpenStudy (anonymous):

u know how hoblos to do this? What is the 7th term of the geometric sequence where a1 = -2 and a5 = -512?

OpenStudy (hoblos):

a5 = a1*r^4 r^4 = a5/a1 = 512/2 = 256 so r = fourth-root(256) = 4 a7 = a1*r^6 = (-2)*(4)^6 = -2(4096) = -8192